find the linear depreciation equation for the TV in the advertisement shown below:
Three (3) years old flat screen TV;
Matrix surround sound with PIP remote
Bought new-$7,500.00; Asking $2,500.00

Answer by AnlytcPhil(1810)

find the linear depreciation equation for the TV in the 
advertisement shown below: 
Three (3) years old flat screen TV;
Matrix surround sound with PIP remote
Bought new-$7,500.00; Asking $2,500.00

----------------------------------------

When it was bought new, the time was = 0 years and the 
value was $7500.

Now that it is being sold, the time is = 3 and the value 
is $2500.

This amounts to finding the equation of a line that 
passes through the two points (0, 7500) and (3, 2500), 
where x represents the time and y represents the value of 
the TV.

     y2 - y1
m = ---------  
     x2 - x1

where (x1,y1) = (0,7500) and (x2,y2) = (3,2500)

     2500 - 7500     -5000
m = ------------- = -------
        3 - 0          3

Now use the point-slope formula:

  y - y1 = m(x - x1)

y - 7500 = -5000/3(x - 0)

y - 7500 = (-5000/3)x

       y = 7500 - (5000/3)x

We now would probably change y to V for value, and
x to t for time

       V = 7500 - (5000/3)t  where 0 < t < 4.5 

Edwin